OpenStudy (anonymous):

Can soomeone help me calculate the bases of matrix 2 0 0 -6 0 2 0 -3 -3 -3 -1 -9 0 0 0 -1 with lamda = -1 & -2

OpenStudy (anonymous):

Do you mean find the eigenvectors?

OpenStudy (anonymous):

Well i need to diagonalize the matrix, I have the answer, but i don't know how to get it

OpenStudy (anonymous):

Did you already find the eigenvalues to be -1 and -2?

OpenStudy (anonymous):

The answers for lamda =-1 v1= 2 1 0 1

OpenStudy (anonymous):

and v2= 0 0 1 0

OpenStudy (anonymous):

no its positive 2

OpenStudy (anonymous):

And for lamda = 2 1 -1 0 0 and 0 1 -1 0

OpenStudy (anonymous):

The problem is i have no idea how to get the answer when i have lines with 0's

OpenStudy (anonymous):

So then your D=(P^-1)A P Where P is the matrix with those vectors as columns. Your D

OpenStudy (anonymous):

i

OpenStudy (anonymous):

your D will be the corresponding eigenvalues on the diagonal. (corresponding to which vector you put where in P)

OpenStudy (anonymous):

righ like -1 000 0-100 0020 0002

OpenStudy (anonymous):

my p is equal to 2 0 1 0 1 0 -1 1 0 1 0 -1 1 0 0 0

OpenStudy (anonymous):

the 2 last colonms, after i apply the lamda = 2 I get a reduced matrix of 1 1 1 0 0 0 0 1 0000 0000

OpenStudy (anonymous):

The answer is 1 -1 0 0

OpenStudy (anonymous):

and 0 1 -1 0

OpenStudy (anonymous):

I have no idea how to get those 2 last vectors

OpenStudy (anonymous):

I'm confused as to where you are having difficulty. Were you able to find a basis for the eigenvalue of -1?

OpenStudy (anonymous):

The answsers are there, I just don't know how to find the base of 2

OpenStudy (anonymous):

The method used to get the 2 bases of the matrix

OpenStudy (anonymous):

yea by lamda i meant eigenvectors sorry

OpenStudy (anonymous):

ok lets work on that

OpenStudy (anonymous):

the bases of -1 were 2 1 0 1 and 0 0 1 0

OpenStudy (anonymous):

So for lamda =2 our matrix is: 0 0 0 6 0 0 0 3 3 3 3 9 0 0 0 3 Right?

OpenStudy (anonymous):

Thats right

OpenStudy (anonymous):

I reduce them to 1 1 1 0 0 0 0 1 0 0 0 0 0 0 0 0

OpenStudy (anonymous):

From there, When theres more than one base to find i'm not sure what to do

OpenStudy (anonymous):

Oh ok so we have leading 1s in which columns?

OpenStudy (anonymous):

first and second?

OpenStudy (anonymous):

No those are the rows with leading columns.

OpenStudy (anonymous):

I mean those are the rows with leading ones.

OpenStudy (anonymous):

Columns 1 and 4 have leading ones.

OpenStudy (anonymous):

So far i follow

OpenStudy (anonymous):

So my base is constructed by x1 and x4 ?

OpenStudy (anonymous):

This tells us that x4=0 x3=? so lets call it "t" x2=? so lets call it "s" x1+x2+x3=0

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

So x1=-s -t x2= s x3= t x4 = 0

OpenStudy (anonymous):

xn=(-1,1,0,0)s+(-1,0,1,0)t

OpenStudy (anonymous):

Thats where I don't know what to do

OpenStudy (anonymous):

\[\left[\begin{matrix}x1 \\ x2\\ x3\\x4\end{matrix}\right]=\left[\begin{matrix}-s -t \\ s\\ t\\0\end{matrix}\right]=\left[\begin{matrix}-s -t \\ s+0\\0+ t\\0+0\end{matrix}\right]=\left[\begin{matrix}-1 \\ 1\\ 0\\0\end{matrix}\right]s+\left[\begin{matrix}-1 \\ 0\\ 1\\0\end{matrix}\right]t\]

OpenStudy (anonymous):

The answer I was given is 1 -1 0 0 and 0 1 -1 0

OpenStudy (anonymous):

Well our first element of the basis matches because that is just (-1) times what I put. I don't know about that 2nd one. Where did you get the answer?

OpenStudy (anonymous):

In a lot of different problems , I get to the step you wrote, but that isn't the answer according to my solutions

OpenStudy (anonymous):

Its the sollutions from past exams .. I could send you the pdfs but Its in french

OpenStudy (anonymous):

Its the first question

OpenStudy (anonymous):

Both vectors are correct. That is just a different basis for the same space.

OpenStudy (anonymous):

Do you know how the vectors were obtained? Because if I do my verifications i don't gt the same answers

OpenStudy (anonymous):

If you subtract our v2 by our v1 you will get their v2.

OpenStudy (perl):

it is in french!>>>???

OpenStudy (anonymous):

Yes

OpenStudy (perl):

you know how to speak french?

OpenStudy (anonymous):

I do

OpenStudy (perl):

download an english linear algebra

OpenStudy (anonymous):

you can check your bases elements by multiplying them by the matrix we row reduced to find them. 0 0 0 6 0 0 0 3 3 3 3 9 0 0 0 3

OpenStudy (anonymous):

But are those vectors linearly independant ?

OpenStudy (anonymous):

the product should be the zero vector

OpenStudy (anonymous):

They are always linearly independent if they are the solutions of a homogeneous system.

OpenStudy (anonymous):

Also the eigenvector bases of distinct eigenvalues are orthoganal to each other, but the elements within each bases may not be (but they will be lin indep).

OpenStudy (anonymous):

I'll verify them right away

OpenStudy (anonymous):

I verified the ones for lamda = 2 already. We are correct.

OpenStudy (anonymous):

Your textbook might benefit from a disclaimer "answer not unique".

OpenStudy (anonymous):

I verified them , and they check out, but in a lot of cases, the answers are written in the the form that i wrote up, and not the ones that you showed me. Even though the answers check out, to be sure to get the aforementioned vectors, we do v2-v1 and what else for the v1? change the signs?

OpenStudy (anonymous):

You don't need to get the same as the textbook. I don't understand your latest question?

OpenStudy (anonymous):

To get v2, the textbook answer, we did v2-v1 to get the v1 answer, we do v1=-v1?

OpenStudy (anonymous):

There is nothing special about the textbook answer. There are infinite bases to describe any vector space (all the linear combinations of the elements in them). If your bases satisfies the multiplication by the matrix = zero vector that is the quickest way to check. Otherwise you are going to have to find a linear combination of your bases that equals each texbook element. That is inefficient.

OpenStudy (anonymous):

But the thing is that those pproblems were corrections from previous exams , and orur teacher wrote it that way, ... im guessing as long as \[P ^{-1}AP\] checks out theresholdnt be any problem right ?

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Yes if that = a diagonal matrix then you are golden.